OpenStudy (anonymous):

Solve Y/-2+ 5 = −4 2 −2 18 −18

OpenStudy (anonymous):

(10-Y)/2 = -4

OpenStudy (anonymous):

Y = 18

OpenStudy (anonymous):

Ok,, Explain

OpenStudy (anonymous):

10-y=-4*2 -y=-8-10 -y=-18 so negative=negative is positive so the ans is 18

OpenStudy (anonymous):

@Deldrickg did u understand this